Support Services Meeting Notes
January 7, 2008
January 7, 2008
Present: Bill Covey, Will Chaney, Kelley Cunningham, Brian Keith, Erich Kesse, Barbara Oliver, Rob RobertsDLC
- Staff are returning from the holiday.
- SAN is up - we'll start using it directly as soon as we are sure replication is working. Will thinks that it is functioning.
- Staff are back from being sick, but now Pete has been called in for jury duty.
- Project status - tables added to the sixth floor - need to get electricity to them. Go ahead with it - order the supplies.
- Second and third floor hold magnets ordered. The fire contacts last week did not work.
- Second floor restroom hallway project - PO is underway.
- Preservation sink install - PO underway.
- Security conversions to resume this week on the problem doors in MSL and Education.
- Preliminary decision on mail conversion - need to discuss mailboxes. UF Central Mail to pre-sort and deliver our mail by department. All vendor addresses will have to be changed. PO Box numbers need to be approved. Will present the proposal to departments in the next couple of weeks and get it approved and in place for campus mail.
- Additionally, we want to get out of the courier service business in the future.
- Current problems - temperature problems in West - there was a flood in the air handling coil, hot water was not available and the air dampers were not working - causing a pipe to break. There were failures on multiple levels. They will need to test it by changing the freezing temperature. The guy should be here to check it out today. The problem seems to be in the water temp handler in the HVAC system.
- Also, there were other AC alarms going off on different floors. Skip is here to check on it.
- Terrazzo in East - go ahead with the smaller projects.
- The Web server is currently up on the new virtual machine. We have good backups for last night. Parts have been ordered for the old Web server. Fix old or leave on new? It would be good if the old server was fixed as the new one is tested.
- Rob is having trouble accessing his files on NT2R1. His folders should be pointing to NT2R, not 2R1.
- Forms-to-mail problem has been fixed.
- Winston is checking on links on the virtual Web server. Shares were not created. Winston is doing it. We paid the company to have replistor create the shares. We also have some licensing issues. It requires one license per virtual server. We will need a handful more of licenses. Will is going to talk to the rep today.
- Brian asks if the PO Box number changes could possibly coincide with the new email address change? No. We may not even have to change our email address with the conversion to campus AD.
- Trudi's last day is January 23. Her retirement tea party will be held January 22 at 3:30PM at the Keene Center. We are working out coverage for her duties after she leaves.
- Dina Benson accepted Gus' old position in DLC. She will start on January 19.
- Positions in Education and ILL were filled.
- We are currently advertising the post of Director of Development.
- Two other Library positions are posted.
- Sam Huang to start no later than March 1st. The Development Officer will report to him.
- We're busy closing out the second quarter.
- The Directors have been scarce in number over the holidays.
- FCLA has determined the size of the grant for the Libraries. It is unknown how it will be split. This is legislative money from FCLA. The amount is not reduced this year. $240,000 has already been corralled by Business Services for Library purchases.
- All of the Directors will soon be running their own blogs. They are expecting them to be commented upon. Startup will initially be through Yahoo's Blogger. Our internal blogger is not being used in production yet. Success of this venture to be determined.
- Division reports - Bill is requesting a departmental report of 2-3 paragraphs encapsulating what was great in your department in the fiscal year of '06-'07. Please get this done at first opportunity. Definitely by the end of the week.